Should print the word Release and the current date \today (in english) on the first two pages, light, outlined. Should print the word Release and current date \today (in english) on bottom of the page after the second page. Check that the lines with ˜ have correct spacing.

\usepackage[light,firsttwo,outline,bottomafter]{draftcopy}

and

\draftcopyName{Release\space\space\space\today}{130}
\draftcopySetScale{80}

in the preamble Test9: This is a test page 1 test the tilde tilde tilde (1 fixed space (˜ ) between the words) test the tilde tilde tilde (2 fixed spaces (˜ ) between the words) test the tilde tilde tilde (3 fixed spaces (˜ ) between the words)

This is a testpage 1

Test9: This is a test page 2 test the tilde tilde tilde (1 fixed space (˜ ) between the words) test the tilde tilde tilde (2 fixed spaces (˜ ) between the words) test the tilde tilde tilde (3 fixed spaces (˜ ) between the words)

This is a testpage 2

Test9: This is a test page 3 test the tilde tilde tilde (1 fixed space (˜ ) between the words) test the tilde tilde tilde (2 fixed spaces (˜ ) between the words) test the tilde tilde tilde (3 fixed spaces (˜ ) between the words)

This is a testpage 3

Test9: This is a test page 4 test the tilde tilde tilde (1 fixed space (˜ ) between the words) test the tilde tilde tilde (2 fixed spaces (˜ ) between the words) test the tilde tilde tilde (3 fixed spaces (˜ ) between the words)

This is a testpage 4